Assume the random variable X has a binomial distribution with the given probability of obtaining a success. Find the following probability, given the number of trials and the probability of obtaining a success. Round your answer to four decimal places.

P(X=14), n=19, p=0.6
Answer #1

Solution

Given that ,

p = 0.6

q = 1 - p = 1 - 0.6=0.4

n = 19

Using binomial probability formula ,

P(X = x) = (n C x) * p x * (1 - p)n - x

P(X = 14) = (19 C 14) * 0.6 14 * (0.4)19 - 14

PROBABILITY=0.0933